# Mesoscopic superfluid to superconductor transition

### v1

@@ VIII.4 Entanglement
Figure 7: Multi-stability versus mode frequency. The diagrams show the dependence of the color-coded spectrum on the mode frequency $\omega_{0}$. The 3 panels from top to bottom are: (a) RGB-coded occupation; (b) The normalized occupation $n_{0}$ of the zero momentum orbital; (c) The occupation $n_{osc}$ of the cavity mode. The other parameters are $\alpha=2$ and $u=3$, while $N=12$ and $N_{osc}=60$. As $\omega_{0}$ is increased the low states go through an SC region (where we have multi-stability); and an SF regime (where we have single-stability). Above the SC/SF region we have a fragmented regime (FR). To the right of the SC/SF/FR regimes we have an $n_{0}$ rainbow region where $n_{osc}=0,1,2$.
@@ VIII.5 Borders
The dependence of the spectrum on $\omega_{0}$ is shown in Fig. 7, featuring SC/SF/FR regions. It is further discussed in the next section.
@@
IX The SC/SF/FR spectral regions
The tomography allows the identification of SC/SF/FR/MI regions in the full quantum spectrum. As model parameters are varied, it is not just the ground state that is affected. We already identified in previous section the MI region using (u,E){(u,E)} diagrams. We now focus on the identification of the SC/SF/FR regions, keeping away from the MI region (i.e. we assume small $u$). The most illuminating figure is possibly Fig. 7b with complementary panels (a, c). In order to understand the illustrated dependence of the spectrum on $\omega_{0}$ we have to clearly distinguish between classical and quantum borders. The classical borders are determined by the classically-scaled parameters $(NU/K,\ \omega_{0}/K,\ N\alpha)$}, whereas quantum borders involve extra $N$ dependence ($1/N$ plays the role of Planck constant). For example the MI border is $u_{L}\sim N^{2}$} for the ground state and $u_{L}\sim N$} for excited states as discussed in [18].
The semiclassical analysis of meta-stability has implicitly assumed that the topography of the energy floor, and in particular of the Landau grooves, can be resolved quantum mechanically. The quantum borders are implied by inspection of the conditions for that. From the perspective of Eq. (14), the question is whether or how many eigenstates can be accommodated by a given valley.
As a preliminary stage let us discuss the quantum significance of $\alpha$. If $\alpha<1$} there is a sequence of $\sim 1/\alpha$} lowest states that have energy spacing $\sim\omega_{0}$, and are contained in the $k_{0}$ groove. The next states belong to the $k_{1}$ groove whose energy floor is higher, namely $W_{\Phi}=\omega_{0}/\alpha$. From now on let us assume for simplicity that $\alpha>1$}, meaning that subsequent states belong to different grooves. Let us call it “scrambled SC region” of the spectrum.
The next question that arises is how many eigenstates can be accommodated by each groove. The rough estimate is determined by the inequality
@@ IX The SC/SF/FR spectral regions
This determines the scrambled SC region in Fig. 7b. This region features mix of blue and red eigenstates, indicating multi-stability. The $1/\omega_{0}$ dependence of its border is apparent. As $\omega_{0}$ is further increased the classical SC-SF border is crossed, and then only the blue eigenstates remain. Above the SC/SF region we see yellowish eigenstates, indicating the FR region.
The right-most quantum region that we see in Fig. 7b is $\omega_{0}>NK$}. This condition implies that the eigenstates do not resolve the topography of the potential floor. The orbitals form an energy band of width $2K$, and therefore all the possible manybody excitations of the ring are in the range whose width is $2NK$. Consequently, the spectrum separates into sequences of states that have the same $\nu_{osc}$, as confirmed by Fig. 7c.
@@ XI Summary and discussion
Formally, in the absence of an EM mode, $\Phi$ can be interpreted as the scaled rotation velocity (Sagnac phase) of the ring. Experiments with such configurations have been realized in cold atom atomtronic experiments [28]. The optional interpretation of $\Phi$ as a dynamical variable that represents an EM mode, requires further attention [29]. Condensation of charged bosons is not a trivial matter [30, 31]. The idea actually predates BCS theory [32], and might be relevant to fermionic systems that feature a BCS-BEC crossover [33].
@@
We have presented a semiclassical tomographic approach to study the SF-SC-FR-MI regions in the $(U,\alpha,\omega_{0},E)$} regime diagram of the circuit. Roughly speaking there are two dimensionless classical parameters $u_{L}$ and $w_{L}$ that control the emergence of SF and CD respectively, and an additional dimensionless quantum parameter $\gamma_{L}$ that controls the MI transition. Illustrations that demonstrate the different spectral regions were provided, notably Fig. 7b and Fig. 8c. The observed spectral regions reflect the underlying valley structure of the energy landscape. The latter features Landau grooves that can support meta-stable condensates.

### v2

@@ VIII.4 Entanglement
Figure 8: Multi-stability versus mode frequency. The diagrams show the dependence of the color-coded spectrum on the mode frequency $\omega_{0}$. The 3 panels from top to bottom are: (a) RGB-coded occupation; (b) The normalized occupation $n_{0}$ of the zero momentum orbital; (c) The occupation $n_{osc}$ of the cavity mode. The other parameters are $\alpha=2$ and $u=3$, while $N=12$ and $N_{osc}=60$. As $\omega_{0}$ is increased the low states go through an SC region (where we have multi-stability); and an SF regime (where we have single-stability). Above the SC-SF region we have a fragmented regime (FR). To the right of the SC-SF-FR regimes we have an $n_{0}$ rainbow region where $n_{osc}=0,1,2$.
@@ VIII.5 Borders
The dependence of the spectrum on $\omega_{0}$ is shown in Fig. 8, featuring SC-SF-FR regions. It is further discussed in the next section.
@@
IX The SC-SF-FR spectral regions
The tomography allows the identification of SC-SF-FR-MI regions in the full quantum spectrum. As model parameters are varied, it is not just the ground state that is affected. We already identified in previous section the MI region using (u,E){(u,E)} diagrams. We now focus on the identification of the SC-SF-FR regions, keeping away from the MI region (i.e. we assume small $u$). The most illuminating figure is possibly Fig. 8b with complementary panels (a, c). In order to understand the illustrated dependence of the spectrum on $\omega_{0}$ we have to clearly distinguish between classical and quantum borders. The classical borders are determined by the classically-scaled parameters $(NU/K,\ \omega_{0}/K,\ N\alpha)$}, whereas quantum borders involve extra $N$ dependence ($1/N$ plays the role of Planck constant). For example the MI border is $u_{L}\sim N^{2}$} for the ground state and $u_{L}\sim N$} for excited states as discussed in [18].
The semiclassical analysis of meta-stability has implicitly assumed that the topography of the energy floor, and in particular of the valleys and the grooves, can be resolved quantum mechanically. The quantum borders are implied by inspection of the conditions for that. From the perspective of Eq. (14), the question is whether or how many eigenstates can be accommodated by a given valley.
As a preliminary stage let us discuss the quantum significance of $\alpha$. If $\alpha<1$} there is a sequence of $\sim 1/\alpha$} lowest states that have energy spacing $\sim\omega_{0}$, and are contained in the $k_{0}$ valley. The next states belong to the $k_{1}$ valley whose energy floor is higher, namely $W_{\Phi}=\omega_{0}/\alpha$. From now on let us assume for simplicity that $\alpha>1$}, meaning that subsequent states belong to different valleys. Below we call it the “scrambled SC region” of the spectrum.
In the tomographic image of the spectrum we can count how many eigenstates we have in each valley if the RGB color-code is used. In a scrambled region e.g. panel (b3) of Fig. 5, we see a mix of the 3 colors. In Fig. 8b the color-code discriminates the $k_{0}$ condensate (blue) from the $k_{\pm}$ condensates (red) because it reflects the occupation of the zero-momentum orbital. Accordingly, the scrambled region features mix of blue and red eigenstates, indicating multi-stability.
The question arises, how many eigenstates can be accommodated by each valley. The rough estimate is determined by the inequality
@@ IX The SC-SF-FR spectral regions
The border of the scrambled SC region in Fig. 8b is determined by Eq. (29). More precisely it should be multiplied by the number of valleys that are being filled (3 in our numerical example). The $1/\omega_{0}$ dependence of its border is apparent.
We continue the discussion of regimes that appear in Fig. 8b. As $\omega_{0}$ is further increased the classical SC-SF border is crossed, and then only the blue eigenstates remain. Above the SC-SF region we see yellowish eigenstates, indicating the FR region. The color-code indicates that the occupation is fragmented as opposed to the condensation that is implied or hinted by either blue or red color respectively.
Finally, let us discuss the right-most quantum ‘rainbow’ region that we see in Fig. 8b. This is the region where $\omega_{0}$ is so large, such that the oscillator is ‘frozen’. Within the lowest rainbow, all the eigenstates feature $\nu_{osc}=0$} (vacuum state), and zero-point fluctuations around the minimum ($\phi\sim n_{osc}\sim 0$}). The color-code merely distinguishes different disentangled eigenstates of a ring that experiences $\phi\sim 0$}. The blue color of the lowest eigenstate in the rainbow reflects condensation in the zero momentum orbital. Assuming negligible $u$, the orbitals form an energy band of width $2K$, and therefore all the possible manybody excitations of the ring are in the range whose width is $2NK$, (or somewhat larger due to the interaction term). Accordingly, the rainbow regime appears in the range $\omega_{0}>NK$}. Consequently, the spectrum separates into sequences of states that have the same $\nu_{osc}$, as confirmed by Fig. 8c.
@@ XI Summary and discussion
Formally, in the absence of an EM mode, $\Phi$ can be interpreted as the scaled rotation velocity (Sagnac phase) of the ring. Experiments with such configurations have been realized in cold atom atomtronic experiments [28]. The alternative interpretation of $\Phi$ as a dynamical variable that represents an EM mode, requires further attention [29]. Condensation of charged bosons is not a trivial matter [30, 31]. The idea actually predates BCS theory [32], and might be relevant to fermionic systems that feature a BCS-BEC crossover [33].
@@
We have presented a semiclassical tomographic approach to study the SF-SC-FR-MI regions in the $(U,\alpha,\omega_{0},E)$} regime diagram of the circuit. Roughly speaking there are two dimensionless classical parameters $u_{L}$ and $w_{L}$ that control the emergence of SF and CD respectively, and an additional dimensionless quantum parameter $\gamma_{L}$ that controls the MI transition. Illustrations that demonstrate the different spectral regions were provided, notably Fig. 8b and Fig. 9c. The observed spectral regions reflect the underlying valley structure of the energy landscape. The latter features valleys and grooves that can support meta-stable condensates.